Suppose that $2000 is deposited in an account that earns 6% interest. How much is in the account
a. after 5 years?
b. after 10 years?
c. after 20 years?
d. after 50 years?
e. after 100 years?
1
Expert's answer
2021-09-09T00:33:17-0400

"a)\\
A= P(1+r)^n= 2000(1+0.06)^5 =\$ 2676.45\\
b)\\
A= P(1+r)^n= 2000(1+0.06)^{10} =\$ 3581.69\\
c)\\
A= P(1+r)^n= 2000(1+0.06)^{20} =\$ 6414.27\\
d)\\
A= P(1+r)^n= 2000(1+0.06)^{50} =\$ 36840.30\\
e)\\
A= P(1+r)^n= 2000(1+0.06)^{100} =\$ 678604.16\\"